# Point-In-Context: Understanding Point Cloud via In-Context Learning

### v1

@@ 5.4.1 Ablations on PIC-G
TABLE VII: Ablation study on PIC-S. We conduct extensive experiments on the Human & Object Segmentation Datasets using our extended version, PIC-S. (a) shows the mIoU gap between using random prompts and ideal prompts to demonstrate the impact of In-Context Labeling (ICL) on the upper limit of PIC-S. Note that AKB-48 is still not included in the training set. (b) shows the effectiveness of In-Context Enhancing (ICE) and In-Context Labeling (ICL). (c) exhibits the comparison results in various choices of the loss function. Gray: default setting.
(a) Influence of In-Context Labeling (ICL)
@@
PIC-S-Cat ($w/o$ ICL) Random 75.15 $\uparrow$ 2.86 82.03 $\uparrow$ 1.74 86.48 $\uparrow$ 0.05 10.30 $\uparrow$ 0.39
@@
PIC-S-Cat ($w/$ ICL) Random 84.13 $\uparrow$ 4.33 80.11 $\uparrow$ 3.04 88.03 $\uparrow$ 0.34 73.52 $\uparrow$ 8.23
@@
PIC-S-Sep ($w/o$ ICL) Random 80.16 $\uparrow$ 0.85 80.85 $\downarrow$ 0.03 84.62 $\downarrow$ 0.05 11.77 $\uparrow$ 0.04
@@
PIC-S-Sep ($w/$ ICL) Random 85.14 $\uparrow$ 2.86 82.82 $\uparrow$ 1.53 88.63 $\uparrow$ 0.26 63.08 $\uparrow$ 14.03
@@
(b) Ablations on In-Context Enhancing  and In-Context Labeling.
  Models ICE ICL
ShapeNetPart
@@
1 $CD$ 84.86 82.60 88.53
2 $Smooth~{}L1$ 84.44 81.88 88.64
3 $CD+Smooth~{}L1$ 85.14 82.82 88.63
@@
Ablation on Point Sampling. We study how the sampling method used by PIC-G in JS modules affects the performance of the model. We use the two most common sampling methods: Farthest Point Sampling (FPS) [38], and Random Sampling (RS) [33]. As shown in Tab. 9(c) (a), for the tasks of reconstruction, denoising, and registration, FPS produces better results than RS. Especially for point cloud reconstruction, where FPS can collect more key points than RS, and these key points can describe the original outline of the entire point cloud. However, in the task of part segmentation, the results of RS exceed those of FPS.
@@
Prompt Engineering. We investigate the influence of the prompt layout and query on the experimental results. For PIC-Sep, we set up two layout options: one with the prompt before the query and the other with the prompt after the query. As shown in Tab. 9(c) (b), the performance between the two designs has neglectable differences. We simply choose the “before" option to align with 2D in-context learning works.
@@
Loss Function. We explore to determine which loss function is most suitable for our PIC-G model. During training, we experiment with using Camfer Distance $\ell_{1}$, $\ell_{2}$, and a combination of $\ell_{1}$ and $\ell_{2}$ as the loss functions for our PIC-Sep. As Tab. 9(c) (c) shows, $\ell_{2}$ achieves the best result on all four tasks.
5.4.2 Ablations on PIC-S
@@
Influence of In-Context Labeling. We conduct an in-depth investigation into the impact of In-Context Labeling (ICL) on our PIC model by comparing its performance with and without the ICL training strategy. We experiment with two prompt selection strategies: random prompts and ideal prompts. The ideal prompt refers to using the query point cloud and its corresponding target point cloud as its prompt, which is technically the most ideal prompt and can lead to the best performance of the model. As shown in Tab. VII (a), by comparing the mIoU distances between using random and ideal prompts, we observe that the model trained with the ICL strategy has a higher upper limit of segmentation performance and greater potential.
@@
Effectiveness of In-Context Enhancing. We also conduct detailed ablation experiments on ICE and ICL. As shown in Tab. VII (b), both ICE and ICL are essential for enhancing the segmentation performance of PIC-S. We observe that using only the ICL training strategy increases the training challenge and may even lead to a decrease in segmentation performance due to introducing a significant amount of randomness. On the other hand, the ICE training strategy, which involves incorporating auxiliary training point cloud pairs, reinforces the model’s learning of mapping relationships in prompts. Therefore, we argue that the best choice is to use ICE and ICL training strategies simultaneously. This can maximize the segmentation performance of the model under multi-dataset joint training.
Loss Function. We conduct ablation experiments on the loss function of PIC-S. As shown in Tab. VII (c), compared with using CD loss or Smooth-$\ell_{1}$ loss alone, the best results can be obtained by combining the two.
@@ 6 Conclusion
We propose Point-In-Context (PIC), the first framework adopting the in-context learning paradigm for 3D point cloud understanding. Specifically, we set up an extensive dataset of point cloud pairs with four fundamental tasks to achieve in-context ability. We propose effective designs that facilitate the training and solve the inherited information leakage problem. PIC shows its excellent learning capacity, achieves comparable results with single-task models, and outperforms multitask models on all four tasks. Besides, it shows good generalization ability to out-of-distribution samples and unseen tasks, and has great potential via selecting higher-quality prompts. We hope it paves the way for further exploration of in-context learning in the 3D modalities.
@@
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. 62203476), and the Natural Science Foundation of Shenzhen (No. JCYJ20230807120801002). This study is also supported under the RIE2020 Industry Alignment Fund Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contributions from the industry partner(s). It is also supported by Singapore MOE AcRF Tier 2 (MOE-T2EP20221-0011). It is also supported by the interdisciplinary doctoral grants (iDoc 2021-360) from the Personalized Health and Related Technologies (PHRT) of the ETH domain.

### v2

@@ 5.4 Ablation Study
Table 9: Ablation study on PIC++. We conduct extensive experiments on the Multi-Entity In-Context Datasets using our extended version, PIC++. (a) shows the mIoU gap between using random prompts and ideal prompts to demonstrate the impact of In-Context Labeling (ICL) on the upper limit of PIC++. Note that AKB-48 is still not included in the training set. (b) shows the effectiveness of In-Context Labeling (ICL) and In-Context Enhancing (ICE). (c) exhibits the comparison results in various choices of the loss function. Gray: default setting.
(a) Influence of In-Context Labeling (ICL) on PIC++
@@
”-Cat” ($w/o$ ICL) Random 75.15 $\uparrow$ 2.86 82.03 $\uparrow$ 1.74 86.48 $\uparrow$ 0.05 10.30 $\uparrow$ 0.39
@@
”-Cat” ($w/$ ICL) Random 84.13 $\uparrow$ 4.33 80.11 $\uparrow$ 3.04 88.03 $\uparrow$ 0.34 73.52 $\uparrow$ 8.23
@@
”-Sep” ($w/o$ ICL) Random 80.16 $\uparrow$ 0.85 80.85 $\downarrow$ 0.03 84.62 $\downarrow$ 0.05 11.77 $\uparrow$ 0.04
@@
”-Sep” ($w/$ ICL) Random 85.14 $\uparrow$ 2.86 82.82 $\uparrow$ 1.53 88.63 $\uparrow$ 0.26 63.08 $\uparrow$ 14.03
@@
(b) Ablations of In-Context Enhancing and In-Context Labeling on the ”-Sep” and ”-Cat” variants of PIC++.
@@
1 CDCD 84.86 82.60 88.53
2 $Smooth-\ell_{1}$ 84.44 81.88 88.64
3 $CD+Smooth-\ell_{1}$ 85.14 82.82 88.63
5.4.1 Ablations on PIC
Effectiveness of Joint Sampling. We investigate whether the JS module has a valid effect on four tasks. As shown in Tab. 5, both PIC-Sep and PIC-Cat face a sharp decline in performance on the four tasks when the JS module is absent, and they even fail to achieve the most basic goal of reconstructing masked tokens. These findings validate our intuition that maintaining the consistency of the input and target token sequence positions is an indispensable design. That is to say, the JS module is a simple yet effective way to compensate for the missing positional information.
Table 10: Ablation experiments with different label point bank sizes $N_{B}$. S., H., B., and A. represent ShapeNetPart, Human3D, BEHAVE, and AKB-48, respectively. We report the ins. mIoU metric.
$N_{B}=8$ $N_{B}=20$ $N_{B}=50$
S. H. B. A. S. H. B. A. S. H. B. A.
PIC++ 85.14 82.82 88.63 63.08 84.37 81.28 88.42 64.47 82.73 70.52 85.89 61.74
Influence of Prompt Selecting Strategies. We explore the impact of the prompt selection on the model’s prediction results. For the random selection method, during testing, we randomly select a pair of input-target point clouds from the training set that performs the same task as the query point cloud, serving as a task prompt. For the class-aware selection method, based on the previous one, we further select point clouds belonging to the same category as the query, such as airplanes, tables, chairs, etc. Additionally, we further delve into selecting two alternative prompts. To select pairs of examples that are paired with the query point cloud, we consider two factors: the Chamfer Distance (CD) [chamferdistance] between the prompt and the query point cloud, and the feature similarity between them (features are extracted from pre-trained PointNet [pointnet]), which are respectively denoted as CD-aware and Fea-aware. As depicted in Tab. 6(b) (a)(b), the CD-aware method demonstrates the best performance and even outperforms the task-specific models on the registration task. Note that we report the random method in the main results (Tab. 2), which means our model has a higher ceiling. This provides us a great opportunity to improve downstream task results by selecting higher-quality prompts, which will be the direction of future work.
[…]
@@ 5.4.1 Ablations on PIC
Ablation on Point Sampling. We study how the sampling method used by PIC in JS modules affects the performance of the model. We use the two most common sampling methods: Farthest Point Sampling (FPS) [pointnet], and Random Sampling (RS) [rs]. As shown in Tab. 9(c) (a), for the tasks of reconstruction, denoising, and registration, FPS produces better results than RS. Especially for point cloud reconstruction, where FPS can collect more key points than RS, and these key points can describe the original outline of the entire point cloud. However, in the task of part segmentation, the results of RS exceed those of FPS.
@@
Prompt Engineering. We investigate the influence of the prompt layout and query on the experimental results. For PIC-Sep, we set up two layout options: one with the prompt before the query and the other with the prompt after the query. As shown in Tab. 9(c) (b), the performance between the two designs has negligible differences. We simply choose the “before” option to align with 2D in-context learning works.
@@
Loss Function. We explore to determine which loss function is most suitable for our PIC model. During training, we experiment with using Chamfer Distance $\ell_{1}$, $\ell_{2}$, and a combination of $\ell_{1}$ and $\ell_{2}$ as the loss functions for our PIC-Sep. As Tab. 9(c) (c) shows, $\ell_{2}$ achieves the best result on all four tasks.
5.4.2 Ablations on PIC++
@@
Influence of In-Context Labeling. We conduct an in-depth investigation into the impact of In-Context Labeling (ICL) on our PIC model by comparing its performance with and without the ICL training strategy. We experiment with two prompt selection strategies: random prompts and ideal prompts. The ideal prompt refers to using the query point cloud and its corresponding target point cloud as its prompt, which is technically the ideal prompt and can lead to the best performance of the model. As shown in Tab. 9(c) (a), by comparing the mIoU distances between using random and ideal prompts, we observe that the model trained with the ICL strategy has a higher upper limit of segmentation performance and greater potential.
@@
Effectiveness of In-Context Labeling and Enhancing. We also conduct detailed ablation experiments on ICL and ICE. As shown in Tab. 9(c) (b), both ICL and ICE are essential for enhancing the segmentation performance of PIC++. We observe that using only the ICL training strategy increases the training challenge and may even lead to a decrease in segmentation performance due to introducing a significant amount of randomness. On the other hand, the ICE training strategy, which involves incorporating auxiliary training point cloud pairs, reinforces the model’s learning of mapping relationships in prompts. Therefore, we argue that the best choice is to use ICE and ICL training strategies simultaneously. This can maximize the segmentation performance of the model under multi-dataset joint training.
Ablation on Different Sizes of Label Point Bank ($N_{B}$). We conducted ablation experiments with different $N_{B}$ values. As shown in Tab. 10, although increasing the bank size to 20 and 50 causes a slight decline in performance, the model maintains competitive results; for instance, the score on ShapeNetPart only decreases marginally from 85.14 to 84.37 ($N_{B}=20$) and 82.73 ($N_{B}=50$), while AKB-48 even sees an improvement to 64.47 at $N_{B}=20$. These results demonstrate that our Dynamic In-Context Labeling strategy ensures that the model learns the underlying mapping rules rather than memorizing specific coordinates, allowing it to generalize effectively to new samples with more parts than those seen during training. Consequently, our method allows for the use of a larger $N_{B}$ in practical applications to guarantee that the number of part labels does not overflow, without compromising performance significantly.
Loss Function. We conduct ablation experiments on the loss function of PIC++. As shown in Tab. 9(c) (c), compared with using CD loss or $Smooth-\ell_{1}$ loss alone, the best results can be obtained by combining the two. In the part segmentation task, $Smooth-\ell_{1}$ is more appropriate for supervised learning of point-level prediction results. Meanwhile, CD loss constrains the prediction of the auxiliary training task (ICE), allowing the model to learn more diverse knowledge about mapping relationships from the prompt pair.
@@ 6 Conclusion
We introduce Point-In-Context (PIC), a foundational framework designed to unify 3D point cloud understanding through in-context learning. To mitigate information leakage in masked point modeling, we propose a Joint Sampling module and establish a comprehensive multitask benchmark. PIC demonstrates robust performance across reconstruction, denoising, registration, and segmentation, showing strong capability in generalization. Crucially, we extend this foundation to PIC++, which incorporates In-Context Labeling and In-Context Enhancing strategies to resolve the limitations of fixed label assignments. These contributions enable seamless scalability across diverse segmentation datasets and facilitate generalization to novel categories, achieving state-of-the-art performance on the proposed multi-entity benchmark. Collectively, PIC and PIC++ pave the way for generalist 3D vision models, marking a substantial step from task-specific tuning toward a scalable, in-context learning paradigm.
@@
7 Limitations
Although the proposed PIC and PIC++ demonstrate strong generalization and versatility, certain limitations remain. First, our experiments are conducted solely on object-level point clouds, leaving scene-level point cloud applications unexplored. Second, our framework is designed to process only 3D point clouds with XYZ coordinates, overlooking other modalities such as color and normal vectors, which restricts its applicability in multimodal scenarios. Finally, while contextual labels significantly enhance the model’s generalization ability, they introduce randomness (random labels) during training, potentially leading to instability. Although our contextual augmentation strategy largely mitigates this issue, occasional randomness may still occur.
8 Future Work
Future Work: First, we plan to extend our PIC framework to large-scale scene-level point clouds, aiming to develop a more comprehensive context learning framework for 3D point clouds. Second, we will incorporate additional modalities to enhance contextual learning and generalization, further exploring the application to broader tasks such as 3D detection and captioning.